OpenStudy (annjewell1998):

Solve the system, using substitution. One point for correct steps, one point for correct ordered pair. x = 4y and 2x + 3y = 44 Please help

OpenStudy (mathwizzard3):

Okay so you just put 4y instead of x

OpenStudy (mathwizzard3):

Like this 2(4y) + 3y = 44

OpenStudy (mathwizzard3):

Than distribute 8y + 3y =44

OpenStudy (mathwizzard3):

Combine like terms 11y = 44

OpenStudy (mathwizzard3):

Divide y = 4

OpenStudy (mathwizzard3):

Voila! Hope that helps!!

OpenStudy (mathwizzard3):

I feel stupid considering your offline but whatever :P

OpenStudy (annjewell1998):

Sorry the internet crashed thank you so much!

OpenStudy (anonymous):

hello need help

OpenStudy (anonymous):



OpenStudy (annjewell1998):

Thanks!

OpenStudy (anonymous):

you are welcome

OpenStudy (anonymous):

well i can't take all of the credit

OpenStudy (anonymous):

Anymore questions

OpenStudy (mathwizzard3):

Can I have a medal, please?

OpenStudy (anonymous):

yeah give the madal to him i can't take any

OpenStudy (anonymous):

*medal

OpenStudy (mathwizzard3):

Thanks you!

OpenStudy (anonymous):

great job guys i am on my way bye

OpenStudy (mathwizzard3):

bye

OpenStudy (annjewell1998):

Thanks guys if you're not busy i do have another question

OpenStudy (anonymous):

ok i will help

OpenStudy (annjewell1998):

Solve the system, using substitution. One point for correct steps, one point for correct ordered pair, (a,b). a + b = 4 and 2a + 3b = 11